# An Automatic Graph Construction Framework based on Large Language Models for Recommendation

### v1

@@ Appendix D Implementation details
We adopt AdamW (Loshchilov and Hutter, 2017) as the optimizer. We optimize the backbone models with the learning rate chosen from {1e-3, 2e-3, 1e-2, 2e-2}, the training batch size from {64, 128, 256} and weight decay from {1e-3, 1e-2}. The training processes have a maximum epoch of 100 with a early stop patience of 3 epochs. Other model-specific hyperparameters (e.g., the number of GRU layers, the number of attention heads) are also determined through grid search to achieve the best results. All the experiments are conducted on V100 GPUs.
@@ D.3. Baseline Methods
The time is computed on a single V100 GPU. For CCGNN, we choose a NER model based on BERT-base (Devlin et al., 2018). For TopoI2I and our AutoGraph, we experiment on Vicuna-7b-v1.5 (Chiang et al., 2023). We use vLLM (Kwon et al., 2023) for LLM inference of a batch size 1. No quantization or other inference techniques are used. For AutoGraph, the running time in initial graph construction is the sum of time cost of both user graphs and item graphs, while in incremental node insertion the time cost is averaged for user nodes and item nodes.
Appendix F Case Study

### v2

@@ Appendix D Implementation details
As a model-agnostic framework, we jointly train the GNN weights in our AutoGraph with the downstream backbone recommendation models. The loss function is the common cross-entropy loss for sequential recommendation (Boka et al., 2024). AdamW (Loshchilov and Hutter, 2017) is adopted as the optimizer. We optimize the backbone models with the learning rate chosen from {1e-3, 2e-3, 1e-2, 2e-2}, the training batch size from {64, 128, 256} and weight decay from {1e-3, 1e-2}. The training processes have a maximum epoch of 100 with a early stop patience of 3 epochs. Other model-specific hyperparameters (e.g., the number of GRU layers, the number of attention heads) are also determined through grid search to achieve the best results. All the experiments are conducted on V100 GPUs.
@@ D.3. Baseline Methods
The time is computed on a single V100 GPU. For CCGNN, we choose a NER model based on BERT-base (Devlin et al., 2018). For TopoI2I and our AutoGraph, we experiment on Vicuna-7b-v1.5 (Chiang et al., 2023). We use vLLM (Kwon et al., 2023) for LLM inference of a batch size 1. No quantization or other inference techniques are used. For AutoGraph, the running time in initial graph construction is the sum of time cost of both user graphs and item graphs, while in incremental node insertion the time cost is averaged for user nodes and item nodes. Moreover, the runtime reported in Table 2 includes the processing time following LLM invocation. However, the overall computational cost is dominated by the LLM usage itself, while the additional processing introduces only negligible overhead and thus does not significantly affect the total runtime.
Next, we provide more detailed analysis on the time complexity of our AutoGraph framework. Under the definition of big $O$ notation, our pointwise method does have the $O(N)$ complexity for initial graph construction.
  • Theoretical Analysis. The overall running time for graph construction is given by $Nt_{1}+m$, where $N$ is the number of nodes (i.e., LLM calls), $t_{1}$ is the time required for a single LLM call per node, and $m$ denotes the time for codebook training. Assuming a batch size of $a$ and training epochs of $e$, then $m=\frac{Net_{2}}{a}$, where $t_{2}$ is the time of one forward and backward pass of a training batch for RQ-VAE. Note that big $O$ notation is concerned with the rate of growth of a function, discarding constant multiplicative factors and lower-order terms (Cormen et al., 2022; Sipser, 1996). Hence, we have $O(Nt_{1}+m)=O(N(t_{1}+\frac{et_{2}}{a}))=O(N)$.
  • Empirical Analysis. Even in practice, $t_{1}$ is more than an order of magnitude larger than $\frac{et_{2}}{a}$. For instance, on the ML-1M dataset, $t_{1}\approx 18.50\,\text{s}$, whereas with $e=3$ and $a=512$, $\frac{et_{2}}{a}\approx 0.05\,\text{s}\ll t_{1}$.
Appendix F More experiments
F.1. The strategy for generating semantic factors
Backbone Model Strategy MovieLens-1M Amazon-Books
NG@10 HR@10 MRR GAUC NG@10 HR@10 MRR GAUC
YT-DNN LSH 0.0560 0.1045 0.0544 0.9040 0.0486 0.0956 0.0451 0.8195
HC 0.0578 0.1105 0.0557 0.9103 0.0552 0.1077 0.0501 0.8338
RQ (Ours) 0.0707 0.1221 0.0686 0.9128 0.0609 0.1173 0.0545 0.8362
MIND LSH 0.0469 0.0946 0.0461 0.9053 0.0592 0.1153 0.0516 0.8105
HC 0.0440 0.0861 0.0396 0.9021 0.0574 0.1112 0.0518 0.8106
RQ (Ours) 0.0643 0.1166 0.0622 0.9126 0.0737 0.1373 0.0656 0.8297
F.2. Hyperparameter Study
Backbone Model Levels MovieLens-1M Amazon-Books
NG@10 HR@10 MRR GAUC NG@10 HR@10 MRR GAUC
YT-DNN $N/A$ 0.0426 0.0906 0.0405 0.8963 0.0444 0.0845 0.0410 0.8417
0 0.0572 0.1039 0.0566 0.9080 0.0570 0.1125 0.0511 0.8347
0, 1 0.0707 0.1221 0.0686 0.9128 0.0609 0.1173 0.0545 0.8362
0, 1, 2 0.0678 0.1231 0.0656 0.9101 0.0561 0.1082 0.0510 0.8360
MIND $N/A$ 0.0435 0.0931 0.0410 0.8936 0.0532 0.1029 0.0487 0.8077
0 0.0584 0.1092 0.0566 0.9078 0.0737 0.1373 0.0656 0.8297
0, 1 0.0643 0.1166 0.0622 0.9126 0.0643 0.1203 0.0580 0.8019
0, 1, 2 0.0572 0.1078 0.0554 0.9094 0.0701 0.1355 0.0615 0.8283
In this part, we study the number of codebook levels used for graph construction and show the result in Table 6. Surprisingly, even though we train 3 levels of codebooks for quantization, including all of them for the final graph construction leads to suboptimal results, while using the first two levels generally yields the best performance.
Table 7. The performance of using semantic vectors from different text sources. Original denotes the simple embeddings of original user/item information. LLM denotes the semantic vectors from LLMs. The best result is given in bold.
Dataset Text Source NG@10 HR@10 MRR GAUC
MovieLens-1M Original 0.0398 0.0876 0.0394 0.8921
LLM 0.0707 0.1221 0.0686 0.9128
Amazon-Books Original 0.0438 0.0887 0.0414 0.8001
LLM 0.0609 0.1173 0.0545 0.8362
F.3. The Importance of LLMs in Graph Construction
We provide experiments to evaluate the information gain introduced by LLMs. Specifically, we use a text encoder (i.e., bge-base-en-v1.5 (Xiao et al., 2023)) to embed original item/user descriptions, compared with the LLM-enriched semantic vectors. We leverage YouTubeDNN as the framework, and keep other configurations the same for both types of semantic vectors. The results are shown in Table 7, from which we can see that the original semantic information is unilateral and shallow for recommendation. Actually in practice, we find that the simple embeddings from original texts lead to poor quantization quality, and cause codebook collapse. In contrast, the in-depth semantics from LLMs allow for meaningful graph construction, thus improving recommendation.
F.4. Case Study